OpenStudy (anonymous):

Can someone manipulate 6,4, and 5 to equal 12?

OpenStudy (curry):

how many times can you use the number

OpenStudy (curry):

is multi and division and exponent allowed as well

OpenStudy (anonymous):

anything is allowed and you can use the numbers only once.

myininaya (myininaya):

5+6/2+4=12

OpenStudy (anonymous):

{(6+4)/5}+(6+4)=2+10=12